I NEED HELP!!!
A premium grade coffee sells for $11 a pound. A standard grade coffee sells for $5 a pound. You decide to make an 18 pound mixture of the premium grade and standard grade and sell it for $6 a pound. How many pounds of premium grade would you use? How many pounds of standard grade would you use?

Answer by richwmiller(17219)

You yelled that you need help.
I am sure you do. Are you implying that others don't need help.
Would anyone post a problem if they didn't need help?
Why are you yelling? I just put new batteries in the hearing aide.
Jar Jar has excellent hearing.
p+s=18
11*p+5*s=6*18
p=3 s=15